OpenStudy (anonymous):

What is the solution to the rational equation 4/3-2/x-5=6/3x-15? x = –8 x = –2 x = 2 x = 8

OpenStudy (acacia):

combine the fractions on both sides of the equation and bring x into the numerator

OpenStudy (acacia):

remember you need a common denominator to add or subtract fractions

OpenStudy (acacia):

First write 4/3 + 2/x+5 and -6/3x+15 as a single fraction

OpenStudy (nurali):

LHS: 4/3 + 2/(2-5) = 4/3 + 2/-3 = 2/3 RHS: -6/(3(2) - 15) = -6/-9 = 2/3 so x=2

OpenStudy (acacia):

sorry I probably didn't articulate that well enough write 4/3 + 2/x+5 as a single fraction and -6/3x+15 as a single fraction keep them seperate 6/ 3x+15 = -6/ 3x+15

OpenStudy (anonymous):

Draw the equation for us it be easier.

OpenStudy (anonymous):

Thanks guys(:

OpenStudy (nurali):

Anytime.